Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

married couples

Expert replies
by [email protected] » Mon Apr 27, 2009 7:03 pm
If 4 people are selected from a group of 6 married couples, what is the probability that none of them would be married to each other?


[spoiler]16/33[/spoiler]
Join the discussion
Source: — Problem Solving |

by cramya » Mon Apr 27, 2009 7:44 pm
Picking 4 people out of 12 people(order does not matter since its the same 4 people no matter when u pick whom)

12C4 (Total possible outcomes)

Lets take 1 married couple and see how many different groupings of 4 can be formed including them

2c2 (no of ways of picking 1 couple) * 10c2 (picking 2 people out of 10 people-> In this 2 we could pick another married couple and this does not matter since in this grouping count we are trying to find where there is atleast 1 married couple)

Since there are 6 marrued couples there will be 6 * [ 2c2*10c2] ways

The no of groupings with no married couples( No of favorable outcomes)
= 12c4 - 6 * [ 2c2*10c2]


So the probability we need = No of favorable outcomes / Total possible outcomes

= 12c4 - 6 * [ 2c2*10c2] / 12c4

= 15/33

Must be missing something.....

Regards,
CR
Join the discussion

by lilu » Mon Apr 27, 2009 8:22 pm
We can solve this problem this way:
12/12*10/11*8/10*6/9=16/33

The explanation to this kind of problem is well described here:
https://www.beatthegmat.com/deck-of-cards-t28564.html
Look for shulapa's explanation.
The more you look, the more you see.
Join the discussion

by dumb.doofus » Mon Apr 27, 2009 8:26 pm
There's a simpler way of doing this..

Let the couples be: Aa, Bb, Cc, Dd, Ee, Ff

Based upon the condition given that none of the person selected may be married to the any of the other three that are selected, implies
Number of way of selecting the first person: 12
Number of way of selecting the second person: 10
Number of way of selecting the third person: 8
Number of way of selecting the Fourth person: 6

So total number of ways of selecting these people who are not married to each other = 12x10x8x6 --- (1)

Total number of ways of selecting 4 people out of 12 = 12x11x10x9 ---(2)

so Probability = (1)/(2)
= 12x10x8x6/(12x11x10x9)
= 16/33

that's the answer :-)
One love, one blood, one life. You got to do what you should.
https://dreambigdreamhigh.blocked/
https://gmattoughies.blocked/
Join the discussion

by dumb.doofus » Mon Apr 27, 2009 8:29 pm
Ooops.. I am sorry.. lilu had not replied by the time I was posting my solution.. please do not consider my solution as plagiarizing.
thnx..
One love, one blood, one life. You got to do what you should.
https://dreambigdreamhigh.blocked/
https://gmattoughies.blocked/
Join the discussion

by cramya » Mon Apr 27, 2009 8:48 pm
Thanks Guys!

Shulapa's method is very good & the way to go. I think my explanations also make it look a lot longer than it is....

Anyway coming back to how I did:

I counted some grouptings twice so the right solution based on what I did would be

Picking 4 people out of 12 people(order does not matter since its the same 4 people no matter when u pick whom)

12C4 (Total possible outcomes) = 495

Lets take 1 married couple and see how many different groupings of 4 can be formed including them

2c2 (no of ways of picking 1 couple) * 10c2

Since there are 6 married couples there will be 6 * [ 2c2*10c2] = 270

I must subtract 15 since when there are 2 couples in a group I would have double counted

270-15 = 255

The no of groupings with no married couples( No of favorable outcomes)
= 495 - 255 = 240


So the probability we need = No of favorable outcomes / Total possible outcomes

= 240/ 495
= 16/33



Regards,
CR
Join the discussion

by LalaB » Sat Dec 17, 2011 7:15 am
total=12C4
none of them are married to each other =6C4*2C1*2C1*2C1*2C1=6C4*(2C1)^4 (here we have chosen 4 couples out of 6, and then we got 1 person from each of these 4 couples)

P= (6C4*(2C1)^4)/12C4=16/33
Join the discussion

by Brent@GMATPrepNow » Sat Dec 17, 2011 7:49 am
[email protected] wrote:If 4 people are selected from a group of 6 married couples, what is the probability that none of them would be married to each other?
[spoiler]16/33[/spoiler]
As with many probability questions, we can also solve this by using probability rules.

P(no couples) = P(choose 1st person AND choose 2nd person such that there are no couples AND choose 3rd person such that there are no couples AND choose 4th person such that there are no couples)

P(no couples) = P(choose 1st person) X P(choose 2nd person such that there are no couples) X P(choose 3rd person such that there are no couples) X P(choose 4th person such that there are no couples)

P(1 choose 1st person) = 12/12 [since we can choose anyone]
P(choose 2nd person such that there are no couples) = 10/11 (11 people remaining, 10 are eligible to be picked)
P(choose 3rd person such that there are no couples) = 8/10 (10 people remaining, 8 are eligible to be picked)
P(choose 4th person such that there are no couples) = 6/9 (9 people remaining, 6 are eligible to be picked)

P(no couples) = (12/12)(10/11)(8/10)(6/9) = 16/33

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by ankush123251 » Mon Dec 19, 2011 4:16 am
Followed this method:
Total number of outcomes = 12C4 =495
No. of favorable outcomes = 6C4 + 6C3 * 3C1 + 6C2 * 4C2 + 6C1 * 5C3 + 6C4(all 4 males + 3 males and 1 female and so on) =240
Thus,
Probability = 240/495 =16/33.
Join the discussion